A flywheel of rotational inertia I = 53kg⋅m2 rotates with angular speed 4.0 rad/s. A tangential force of 6.5 N is applied at a distance of 0.36 m from the center in such a way that the angular speed decreases. How long will it take the wheel to stop?

Result: 100 s


1
Expert's answer
2021-01-24T16:42:48-0500

Let's find the torque that acts on the flywheel to stop it:


"\tau=Fr=-6.5\ N\cdot0.36\ m=-2.34\ N\cdot m."

From the other hand, torque can be written as follows:


"\tau=\alpha I."

Then, we can find the angular decceleration of the flywheel:


"\alpha=\dfrac{\tau}{I}=\dfrac{-2.34\ N\cdot m}{53\ kg\cdot m^2}=-0.04\ \dfrac{rad}{s^2}."

Finally, we can find the time that the flywheel takes to stop:


"\alpha=\dfrac{\omega-\omega_0}{t},""t=\dfrac{0-4\ \dfrac{rad}{s}}{-0.04\ \dfrac{rad}{s^2}}=100\ s."